# BSTcontrol

### v2

@@ 4.2 Distributed Collaborative Machine Learning
Distributed collaborative machine learning (DCML) gains popularity due to its data privacy advantages [196]. Unlike traditional methods, data in DCML is accessed collectively without transferring from administrators to any untrusted parties. DCML enables distributed model training on decentralized data. Currently popular DCML methods include FL, SL and FSL.
@@ 4.2.1 Federated Learning
FL encompasses several approaches tailored to different data distribution scenarios, particularly relevant in wireless communications. Horizontal FL (sample-based FL) applies when clients share the same feature space but have different user samples, enabling collaboration between entities like telecom companies without sharing user data [198, 199]. Vertical FL suits scenarios where clients hold different feature sets for the same users, requiring advanced network optimization to combine features effectively [200]. Federated transfer learning addresses situations where both samples and features differ across clients, leveraging shared representations to enhance model training, especially useful for collaboration between diverse wireless service providers or IoT networks with minimal data overlap [201].
@@
A crucial step in the FL process is model aggregation, where the server combines the updates received from clients. Various techniques enhance this step. For instance, robust federated aggregation employs the geometric median to resist potential malicious interference [202]. Selective model aggregation methods evaluate local data quality or computational capacity to choose which client updates to include, optimizing resource use in environments like vehicular edge computing [203]. Secure aggregation protocols use secure multi-party computation to protect the privacy of individual client gradients during aggregation [204, 205].
FL also offers methods to address challenges like insufficient labeled data on client devices. Federated few-shot learning (FedFSL) enables models to classify new classes using very few labeled samples [206]. Private semi-supervised federated learning leverages abundant unlabeled data alongside scarce labeled data in a privacy-preserving manner [207]. Personalized FedFSL goes further by identifying optimal collaborating clients for specific tasks, learning personalized feature spaces without data disclosure [208]. Techniques like FedAffect update feature extractors using disjoint unlabeled private data to learn diverse representations [209]. For LLMs, methods like AUG-FedPrompt use minimal labeled data initially and augment it by annotating unlabeled data, though potentially incurring high communication costs due to full parameter tuning [210].
@@
Fed-LLM can be deployed during the pre-training stage to customize specific needs of end users and improve performance for specialized tasks. The authors in [186] proposed a framework which integrates centralized public data with distributed private data sources for pre-training. This design tailors the LLM architecture based on pre-training parameter selection and task requirements. Multiple clients initiate pre-training with local data and subsequently perform model sharing. The utilization of diverse computing resources not only enhances model generalization but also maintains data privacy. In the context of wireless communication, this approach allows for robust, efficient, and privacy-preserving model training across various network nodes, improving overall system performance and adaptability.
@@
Fed-LLM can also benefit the fine-tuning process of the large AI models in practical wireless environments. In the target deployment scenarios, multiple clients can use local data for federated collaborative parameter-tuning. The fine-tuned model parameters are then uploaded to a server for aggregation and distribution until the model reaches convergence [211]. To reduce the computation and communication costs of full-model fine-tuning, parameter-efficient fine-tuning methods [212] can be integrated into the Fed-LLM framework to minimize parameter gradient calculations and reduce the number of aggregated parameters. This can achieve a balance between maintaining performance and mitigating the computation and communication overhead, which is essential for efficient and scalable model updates.
@@
Additionally, prompt engineering can take advantage of Fed-LLM. Prompt engineering is the process of guiding a large AI model to produce a desired output. Although large AI models attempt to mimic the behaviors of humans, detailed instructions are still required to create high-quality and relevant output [213]. Standard prompt design often relies on public data, limiting applicability for specialized or personalized tasks. The reuse of general prompts from public datasets may also reduce responsiveness of the models. By leveraging FL in wireless communication networks, prompts can be generated using private data while ensuring privacy protection [214]. This approach improves the generalization ability of the models, enabling the model to handle tasks in specialized fields more proficiently. Moreover, personalized prompts can be created to meet the specific needs of clients, making the AI systems more responsive and contextually aware [210].
@@ 4.2.2 Split Learning
SL is a privacy-enhancing machine learning method that enables multiple nodes to collaboratively train a global model without revealing the original data [187]. SL typically divides a machine learning model into two parts [215]: a client-side sub-model and a server-side sub-model, and deploys them on different nodes. Usually, one node acts as the server, hosting the server-side sub-model, while other nodes act as clients, hosting the client-side sub-models. During model training, the client-side and server-side sub-models jointly complete forward propagation and backpropagation in each iteration. Unlike FL, which only employs data parallelism, SL combines data parallelism and model parallelism. This approach is suitable for scenarios with limited client computing and communication capabilities, such as edge computing and the IoT. Additionally, splitting machine learning models across different nodes in SL helps protect privacy.
@@
Currently, many studies on SL have been conducted. In the field of edge computing, researchers evaluate the learning performance and implementation overhead of SL in real-world IoT scenarios [216]. Besides, SL can also be used as a form of distributed learning to achieve URLLC [217]. Furthermore, In institutional collaboration, SL can be used to collaboratively train health models across multiple medical institutions [218]. In terms of privacy protection, some studies have indicated that the activation layer outputs in SL training may lead to data privacy leaks. To address this, various methods and techniques, such as differential privacy, have been proposed to enhance privacy protection [219].
@@ 4.2.3 Federated Split Learning
Integrating the benefits of both SL and FL [220, 188], FSL has garnered significant research interest, with numerous studies currently exploring its potential. The foundational FSL architecture is depicted in Fig. 13. Research efforts have focused on optimizing FSL, particularly through multi-head SL, which operates without client model synchronization. Experiments confirm this approach is viable and achieves performance comparable to traditional FSL [221]. From a security perspective, FSL offers enhanced robustness against model poisoning attacks, as clients handle only partial models and lower-dimensional data [222]. Studies examining data poisoning attacks further reveal that non-targeted attacks by malicious participants impact global model accuracy more significantly than targeted attacks [223]. Additionally, adaptations for specific domains, such as a mobile FSL method for vehicular networks, have demonstrated improved training speeds, substantially reducing training time compared to conventional FSL without sacrificing model accuracy [224].
@@ 4.3 Over-the-Air Computation
AirComp represents a paradigm shift from conventional communication protocols by integrating computation directly into the communication process [11]. This technique is particularly crucial for future wireless networks aiming to support large AI models distributed across numerous devices. The architecture of AirComp is depicted in Fig. 14. Instead of treating concurrent transmissions as interference to be mitigated, AirComp harnesses the waveform superposition property inherent in multiple access channels to compute desired functions, which are often aggregation functions essential for distributed AI tasks from the data transmitted by multiple devices simultaneously [225, 11]. This compute-when-communicate approach contrasts sharply with the traditional compute-after-communicate strategy, offering significant advantages in spectral efficiency and latency. These benefits are paramount for applications involving the massive data aggregation required by distributed large AI model training or inference [11]. The core principle involves designing transceivers such that the superimposed signal at the receiver directly represents a target function (often a summation, potentially after device side preprocessing and followed by receiver side postprocessing), thereby bypassing the need for individual signal decoding and significantly reducing communication bottlenecks associated with large AI models [225]. While AirComp introduces challenges related to managing aggregation errors caused by channel fading and noise, its potential for efficiency gains in supporting distributed AI has spurred research into various techniques and applications, three prominent examples of which are detailed below.
@@ 4.3.1 AirComp for Federated Edge Learning
One of the most significant applications of AirComp is in the domain of federated edge learning (FEEL), leading to the concept of Air-FEEL [189]. FEEL is a key technique for training large AI models collaboratively across distributed edge devices holding local data, without centralizing sensitive user data. However, in conventional FEEL, the communication overhead associated with frequently exchanging high dimensional model parameters or gradients, characteristic of large AI models between numerous edge devices and a central server constitutes a major bottleneck [226]. Air-FEEL addresses this critical challenge by employing AirComp for one shot aggregation of these local updates over the wireless channel [11]. This approach significantly enhances communication efficiency and reduces training latency, making the distributed training of large models more feasible, compared to methods requiring orthogonal resource allocation or sequential decoding [189, 11]. Two primary implementations exist: Air-FedSGD, which aggregates gradients after each local computation, and Air-FedAvg, which aggregates model parameters after multiple local iterations, both facilitating the efficient update of a global large AI model [189]. Despite the introduction of aggregation errors, sophisticated techniques analyze and optimize Air-FEEL performance, linking aggregation errors (bias and mean squared error) to the learning optimality gap of the large model [227, 228]. Optimized power control strategies further aim to minimize this gap, managing the tradeoff between convergence speed and aggregation errors, thus directly impacting the training efficiency of the large AI model [227, 228]. Furthermore, Air-FEEL inherently enhances privacy, a critical concern when dealing with data used to train large models, by masking individual updates within the aggregated signal [189].
4.3.2 Spatial AirFusion for Integrated Sensing and Edge AI

### v4

@@ 4.2 Distributed Collaborative Machine Learning
Distributed collaborative machine learning (DCML) gains popularity due to its data privacy advantages [kairouz2021advances]. Unlike traditional methods, data in DCML is accessed collectively without transferring from administrators to any untrusted parties. DCML enables distributed model training on decentralized data. Currently popular DCML methods include FL, split learning (SL) and federated split learning (FSL), and the important properties are summarized in Table 7.
Table 7: Comparisons of Distributed Learning Frameworks.
Feature Federated Learning Split Learning Federated Split Learning
Core Idea Parallel local training of the full model, followed by central aggregation of the updates. Sequential training of a vertically split model, where clients and a server collaboratively process the model layer by layer. A hybrid approach combining the parallelism of FL with the low client-side workload of SL.
Convergence Analysis The convergence bound reveals a dependency on data heterogeneity $\Gamma$ and the number of local epochs $E$. The convergence guarantee is dependent on the number of clients $N$ and local steps $K$ in deep learning scenarios. The convergence bound depends on local iterations $\tau$ and data heterogeneity factors such as $\sigma_{n}^{2}$ and $\epsilon^{2}$.
System Delay Model Defined by the slowest client’s processing time:
$\max(T_{\text{comp}}+T_{\text{comm}})$. This is known as the straggler effect.
Defined by the cumulative time of all clients processing in sequence:
$\sum(T_{\text{comp}}+T_{\text{comm}})$. This creates a sequential bottleneck.
A hybrid model defined by parallel client tasks and sequential server tasks:
$\max(T_{\text{client}})+\sum(T_{\text{server}})$.
Client Computation Load The clients must process the entire model architecture, which demands significant local memory and computational resources. The computational burden on clients is minimal, as they only process the initial and computationally lighter layers of the model. Client computation is significantly reduced by processing only the initial model layers, making the architecture suitable for edge devices.
Communication Overhead Content: Full model parameters $S_{\text{model}}$.
Frequency: Once per round, at the aggregation stage.
Content: Smashed data and activations $S_{\text{act}}$.
Frequency: Twice per local step (forward and backward).
Content: Smashed data and activations $S_{\text{act}}$.
Frequency: Twice per local step.
@@ 4.2.1 Federated Learning
FL encompasses several approaches tailored to different data distribution scenarios, particularly relevant in wireless communications. Horizontal FL (sample-based FL) applies when clients share the same feature space but have different user samples, enabling collaboration between entities like telecom companies without sharing user data [yang2020horizontal]. Vertical FL suits scenarios where clients hold different feature sets for the same users, requiring advanced network optimization to combine features effectively [liu2024vertical]. Federated transfer learning addresses situations where both samples and features differ across clients, leveraging shared representations to enhance model training, especially useful for collaboration between diverse wireless service providers or IoT networks with minimal data overlap [liu2020secure].
@@
A crucial step in the FL process is model aggregation, where the server combines the updates received from clients. Various techniques enhance this step. For instance, robust federated aggregation employs the geometric median to resist potential malicious interference [pillutla2022robust]. Selective model aggregation methods evaluate local data quality or computational capacity to choose which client updates to include, optimizing resource use in environments like vehicular edge computing [ye2020federated]. Secure aggregation protocols use secure multi-party computation to protect the privacy of individual client gradients during aggregation [bonawitz2016practical, fereidooni2021safelearn].
To better adapt federated learning for practical wireless challenges such as device heterogeneity and dynamic connectivity, recent research has introduced more specialized architectures [shfMIOT], contributing to the broader vision of ’Federated Intelligence’. One such approach is clustered federated fine-tuning, which groups devices with similar characteristics to handle data heterogeneity and train more customized models. Another paradigm, hierarchical federated fine-tuning, employs a multi-level aggregation structure to effectively balance local model personalization with global generalization. Finally, asynchronous federated fine-tuning enhances network efficiency by allowing devices to submit updates independently, thereby accommodating participants with varying processing speeds and intermittent connectivity.
FL also offers methods to address challenges like insufficient labeled data on client devices. Federated few-shot learning (FedFSL) enables models to classify new classes using very few labeled samples [fan2021federated]. Private semi-supervised federated learning leverages abundant unlabeled data alongside scarce labeled data in a privacy-preserving manner [fan2022private]. Personalized FedFSL goes further by identifying optimal collaborating clients for specific tasks, learning personalized feature spaces without data disclosure [zhao2022personalized]. Techniques like FedAffect update feature extractors using disjoint unlabeled private data to learn diverse representations [shome2021fedaffect]. For LLMs, methods like AUG-FedPrompt use minimal labeled data initially and augment it by annotating unlabeled data, though potentially incurring high communication costs due to full parameter tuning [cai2022aug].
@@
Fed-LLM can be deployed during the pre-training stage to customize specific needs of end users and improve performance for specialized tasks. The authors in [OpenFedLLM] proposed a framework which integrates centralized public data with distributed private data sources for pre-training. This design tailors the LLM architecture based on pre-training parameter selection and task requirements. Multiple clients initiate pre-training with local data and subsequently perform model sharing. The utilization of diverse computing resources not only enhances model generalization but also maintains data privacy. In the context of wireless communication, this approach allows for robust, efficient, and privacy-preserving model training across various network nodes, improving overall system performance and adaptability.
@@
Fed-LLM can also benefit the fine-tuning process of the large AI models in practical wireless environments. In the target deployment scenarios, multiple clients can use local data for federated collaborative parameter-tuning. The fine-tuned model parameters are then uploaded to a server for aggregation and distribution until the model reaches convergence [hilmkil2021scaling]. To reduce the computation and communication costs of full-model fine-tuning, parameter-efficient fine-tuning methods [ding2023parameter] can be integrated into the Fed-LLM framework to minimize parameter gradient calculations and reduce the number of aggregated parameters. This can achieve a balance between maintaining performance and mitigating the computation and communication overhead, which is essential for efficient and scalable model updates.
@@
Additionally, prompt engineering can take advantage of Fed-LLM. Prompt engineering is the process of guiding a large AI model to produce a desired output. Although large AI models attempt to mimic the behaviors of humans, detailed instructions are still required to create high-quality and relevant output [liu2023pre]. Standard prompt design often relies on public data, limiting applicability for specialized or personalized tasks. The reuse of general prompts from public datasets may also reduce responsiveness of the models. By leveraging FL in wireless communication networks, prompts can be generated using private data while ensuring privacy protection [fedprompt]. This approach improves the generalization ability of the models, enabling the model to handle tasks in specialized fields more proficiently. Moreover, personalized prompts can be created to meet the specific needs of clients, making the AI systems more responsive and contextually aware [cai2022aug].
To analyze FL performance, we model its distinct communication round. A typical round begins with the central server transmitting the current global model to the $N$ clients of the network. Each client then performs multiple local training iterations on its private dataset before computing a model update. Subsequently, all $N$ clients transmit their updates back to the server. The round concludes when the server aggregates these $N$ updates to produce a refined global model for the next round.
This iterative process directly impacts learning convergence, particularly in the challenging scenario of non-independent and identically distributed (non-IID) client data. For a smooth and strongly convex objective, the convergence of the FedAvg algorithm [Li2020On] is bounded by :
$\mathbb{E}[F(\mathbf{w}_{T})]-F^{*}\leq\frac{2\kappa}{\gamma+T}\left(\frac{B}{\mu}+2L\|\mathbf{w}_{0}-\mathbf{w}^{*}\|^{2}\right),$ (20)
where $B$ is defined as:
$B=\sum_{k=1}^{N}p_{k}^{2}\sigma_{k}^{2}+6L\Gamma+8(E-1)^{2}G^{2}.$ (21)
In (20), $\mathbb{E}[F(\mathbf{w}_{T})]$ is the expected loss of the global model weights $\mathbf{w}_{T}$ after $T$ rounds, and $F^{*}$ is the optimal loss value. The initial model weights are $\mathbf{w}_{0}$ and the optimal model weights are $\mathbf{w}^{*}$. The constants $\mu$ and $L$ represent the strong convexity and smoothness of the loss function, with their ratio defined as $\kappa=L/\mu$. The term $\gamma$ is a constant defined as $\gamma=\max\{8\kappa,E\}$. For the component $B$, NN is the total number of clients, $p_{k}$ is the data proportion for client $k$, and $\sigma_{k}^{2}$ is the variance of its stochastic gradients. The term $\Gamma$ quantifies the degree of data heterogeneity across clients. Finally, $E$ is the number of local training epochs each client performs, and $G^{2}$ is an upper bound on the squared norm of gradients.
The latency of this communication round is also critical. In a synchronous system, the total delay is constrained by the slowest participant [yang2020delay]. This ‘straggler’ effect is a major challenge when many devices are involved. A representative delay model is:
$T_{FL}=\max_{k\in\{1,...,N\}}\left(E\cdot\frac{D_{k}C_{k}}{f_{k}}\right)+\frac{S_{model}}{B_{dl}}+\frac{S_{model}}{B_{ul}}.$ (22)
Here, $T_{FL}$ represents the total time for one round. The first term models the computation delay, where the $\max$ operator is taken over all $N$ clients. For each client $k$, $E$ is the number of local epochs, $D_{k}$ is its local dataset size, $C_{k}$ is the computational cycles required per data sample, and $f_{k}$ is the CPU frequency of the client. The second and third terms represent communication delay, where $S_{model}$ is the model’s parameter size, and $B_{dl}$ and $B_{ul}$ are respectively the available downlink and uplink bandwidth.
@@ 4.2.2 Split Learning
SL is a privacy-enhancing machine learning method that enables multiple nodes to collaboratively train a global model without revealing the original data [10843784]. SL typically divides a machine learning model into two parts [gupta2018distributed]: a client-side sub-model and a server-side sub-model, and deploys them on different nodes. Usually, one node acts as the server, hosting the server-side sub-model, while other nodes act as clients, hosting the client-side sub-models. During model training, the client-side and server-side sub-models jointly complete forward propagation and backpropagation in each iteration. Unlike FL, which only employs data parallelism, SL combines data parallelism and model parallelism. This approach is suitable for scenarios with limited client computing and communication capabilities, such as edge computing and the IoT. Additionally, splitting machine learning models across different nodes in SL helps protect privacy.
Currently, many studies on SL have been conducted. In the field of edge computing, researchers evaluate the learning performance and implementation overhead of SL in real-world IoT scenarios [yansong2020end]. Besides, SL can also be used as a form of distributed learning to achieve URLLC [park2020extreme]. Furthermore, In institutional collaboration, SL can be used to collaboratively train health models across multiple medical institutions [abuadbba2020can]. In terms of privacy protection, some studies have indicated that the activation layer outputs in SL training may lead to data privacy leaks. To address this, various methods and techniques, such as differential privacy, have been proposed to enhance privacy protection [vepakomma2020nopeek].
The communication scenario in SL is fundamentally sequential. In a given training round, the server interacts with $N$ clients of the network one by one in a defined sequence. The process begins with the first client, which performs its local training by exchanging activation tensors and gradients with the server. After completing its local steps, this client passes its updated model to the second client in the sequence. This relay-based process continues until the final client, client $N$, has finished its training, marking the end of the round.
This sequential interaction profoundly affects the framework’s convergence. For a non-convex objective on non-IID data, the convergence of SL is bounded. As shown in [li2023convergence]:
[…]
In this expression, the left side is the convergence metric, where $\overline{x}^{R}$ is the averaged global model over $R$ total rounds. The first term on the right side represents the initialization error, where $f(x^{0})$ is the initial loss and $f(x^{*})$ is the optimal loss. $N$ is the number of clients, and $K$ is the number of local update steps per client. The terms $\eta_{g}$ and $\eta_{l}$ are the global and local learning rates, respectively. The second and third terms collectively represent the client drift error, which depends on the learning rates, the number of clients and local steps, the Lipschitz constant $L$, the bounded gradient norm $G$, and the gradient variance $\sigma^{2}$. The final term represents the global variance.
The system’s latency is a direct reflection of this sequential communication model. The total time for one complete round is the accumulation of the processing times for every single client. This relationship is captured by the delay model:
$T_{SL}=\sum_{i=1}^{N}\left(K\cdot(T_{fp,i}^{c}+T_{comm,i}^{act}+T_{server}+T_{comm,i}^{grad}+T_{bp,i}^{c})\right).$ (24)
In this equation, $T_{SL}$ is the total delay for one round. For each client $i$, the inner parenthesis represents the time for one local step, consisting of client-side forward $T_{fp,i}^{c}$ and backward $T_{bp,i}^{c}$ propagation, server processing $T_{server}$, and communication of activations $T_{comm,i}^{act}$ and gradients $T_{comm,i}^{grad}$. This single-step time is multiplied by $K$, the number of local steps. The summation operator $\sum$ over all $N$ clients is the key feature, showing that total latency scales linearly with the number of participants.
@@ 4.2.3 Federated Split Learning
Integrating the benefits of both SL and FL [thapa2022splitfed, zhao2024fedsllm], FSL has garnered significant research interest, with numerous studies currently exploring its potential. The foundational FSL architecture is depicted in Fig. 14. Research efforts have focused on optimizing FSL, particularly through multi-head SL, which operates without client model synchronization. Experiments confirm this approach is viable and achieves performance comparable to traditional FSL [joshi2021splitfed]. From a security perspective, FSL offers enhanced robustness against model poisoning attacks, as clients handle only partial models and lower-dimensional data [khan2022security]. Studies examining data poisoning attacks further reveal that non-targeted attacks by malicious participants impact global model accuracy more significantly than targeted attacks [gajbhiye2022data]. Additionally, adaptations for specific domains, such as a mobile FSL method for vehicular networks, have demonstrated improved training speeds, substantially reducing training time compared to conventional FSL without sacrificing model accuracy [moon2023split].
FSL redesigns the communication scenario to introduce parallelism. In an FSL round, a Fed Server coordinates with the network’s $N$ clients, while a Main Server handles the server-side model computation. The round begins with the Fed Server distributing the client-side model $x_{c}$ to all $N$ clients. Each client then performs $\tau$ local iterations. In each iteration, all clients in parallel perform a forward pass and transmit their activation tensors to the Main Server. The Main Server sequentially processes these activations, updates its server-side model $x_{s}$, and returns gradients to each client. After clients complete their backward passes, they update their local client-side models. At the round’s end, all $N$ clients send their updated models to the Fed Server for aggregation.
This architecture’s performance has been rigorously analyzed. For a non-convex objective, the convergence of SFL is bounded. As derived in [han2024convergence]:
[…]
Here, the left side is the convergence metric averaged over $T$ total rounds. On the right side, the first term represents the initialization error, where $f(x^{0})-f^{*}$ is the initial loss gap and $\tau$ is the number of local iterations. The second term represents the error accumulation due to data heterogeneity. $N$ is the total number of clients, $S$ is the smoothness constant, and $\eta^{t}$ is the learning rate at round $t$. The term $a_{n}$ is the weight of client $n$, $\sigma_{n}^{2}$ is its local gradient variance, and $\epsilon^{2}$ bounds the divergence between local and global gradients.
The latency of FSL reflects its hybrid parallel-sequential nature. The per-round delay is determined by the slowest client, combined with the sequential processing at the Main Server. A representative delay model is:
$T_{FSL}=\max_{k\in\{1,...,N\}}\left(\tau\cdot(T_{fp,k}^{c}+T_{comm,k}^{act})\right)+\sum_{k\in\{1,...,N\}}T_{server\_proc,k}+\max_{k\in\{1,...,N\}}\left(\tau\cdot(T_{comm,k}^{grad}+T_{bp,k}^{c})\right).$ (26)
Here, $T_{FSL}$ is the total round time. The first and third terms represent the parallel computation and communication at the clients, limited by the slowest client among all $N$ participants (the $\max$ operator). The middle term, $\sum_{k\in\{1,...,N\}}T_{server\_proc,k}$, reflects that the Main Server still processes the activations from all $N$ clients sequentially, which can remain a bottleneck.
@@ 4.3 Over-the-Air Computation
AirComp represents a paradigm shift from conventional communication protocols by integrating computation directly into the communication process [wang2024over]. Instead of treating concurrent transmissions as interference, AirComp harnesses the inherent waveform superposition property of multiple-access channels to compute desired aggregation functions, such as weighted summations, directly over the air [zhu2021over, wang2024over]. This ‘compute-when-communicate’ approach offers significant advantages in spectral efficiency and latency compared to traditional ‘compute-after-communicate’ strategies.
These benefits are particularly crucial for large AI models, whose distributed operations generate massive data aggregation requirements that can overwhelm conventional networks. By designing transceivers to perform in-network computation, AirComp directly addresses the bottleneck of exchanging high-dimensional gradients in federated fine-tuning, especially when combined with PEFT methods [shfTWC]. Furthermore, it is being adapted to support a wide range of large AI models-centric applications, from enabling scalable federated edge learning (FEEL) [cao2024overview] to executing generalized functions for diverse model architectures [liu2023over] and even efficiently fusing sensor data for large multi-modal perception models [liu2024over].
@@ 4.3.1 AirComp for Federated Edge Learning
One of the most significant applications of AirComp is in the domain of FEEL, leading to the concept of scalable FEEL (Air-FEEL) [cao2024overview]. FEEL is a key technique for training large AI models collaboratively across distributed edge devices holding local data, without centralizing sensitive user data. However, in conventional FEEL, the communication overhead associated with frequently exchanging high dimensional model parameters or gradients, characteristic of large AI models between numerous edge devices and a central server constitutes a major bottleneck [XIAO2024]. Air-FEEL addresses this critical challenge by employing AirComp for one shot aggregation of these local updates over the wireless channel [wang2024over]. This approach significantly enhances communication efficiency and reduces training latency, making the distributed training of large models more feasible, compared to methods requiring orthogonal resource allocation or sequential decoding [cao2024overview, wang2024over]. Two primary implementations exist: Air-FedSGD, which aggregates gradients after each local computation, and Air-FedAvg, which aggregates model parameters after multiple local iterations, both facilitating the efficient update of a global large AI model [cao2024overview]. Despite the introduction of aggregation errors, sophisticated techniques analyze and optimize Air-FEEL performance, linking aggregation errors (bias and mean squared error) to the learning optimality gap of the large model [cao2021optimized, cao2022transmission]. Optimized power control strategies further aim to minimize this gap, managing the tradeoff between convergence speed and aggregation errors, thus directly impacting the training efficiency of the large AI model [cao2021optimized, cao2022transmission]. Furthermore, Air-FEEL inherently enhances privacy, a critical concern when dealing with data used to train large models, by masking individual updates within the aggregated signal [cao2024overview].
Building on this principle, recent works have tailored Air-FEEL specifically for the challenges of fine-tuning large AI models. A key strategy is to combine AirComp with PEFT methods like LoRA. The authors in [shfTWC] propose the AirFL-LORA framework, where instead of transmitting the full, high-dimensional gradients, devices only need to transmit the much smaller low-rank update matrices. This framework designs a multi-input multi-output (MIMO)-OFDM-based AirComp scheme to efficiently aggregate these low-rank adapters over the air. By jointly optimizing the rank of the adapters, resource allocation, and beamforming, the framework minimizes the learning optimality gap, achieving a balance between communication efficiency and model performance that is critical for fine-tuning LLMs on resource-constrained edge devices.
4.3.2 AirComp for Large Multi-Modal Models in ISAC